# What Prompts Don’t Say: Understanding and Managing Underspecification in LLM Prompts

### v1

@@ 3.3 LLMs are more likely to regress on unspecified requirements when updated
Figure 3: Cumulative distribution of accuracy drop. Prompts regress more on unspecified requirements across model updates, with an almost 2x increase compared to specified requirements.
Setups. To analyze model updates, we study six models from gpt-4o and llama-3 model families: Three versions of gpt-4o: gpt-4o-2024-05-13, gpt-4o-2024-08-06, and gpt-4o-2024-11-20 for simulating a hidden drift of model versions, and three versions of llama-3: Llama-3-70B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct, simulating intentional model migration and bigger changes. For each potential update within the same model family, we calculate the accuracy drop for each prompt and requirement.
Results. We found that prompts regress more on unspecified requirements: 5.9% requirements regress more than 20% over model updates when they are unspecified – an almost 2x increase compared to specified requirements (Figure 3). This is true even for small hidden model updates, – e.g., updating from gpt-4o-2024-05-13 to gpt-4o-2024-08-06 makes it 48% less likely to “produce skimmable outputs,” and 14% less likely to “mention customer support information” on average.
Implications. As the requirements are unspecified, their regressions will be much harder to notice despite being even more common. This makes it necessary to regularly evaluate and monitor known unspecified requirements (Section 5). Existing practices of manual inspection (“vibe check”) of a few examples will not be sustainable.
3.4 Identifying unspecified requirements is challenging with existing practices
If underspecified prompts are more unstable, could developers discover relevant task requirements in the first place? We argue that this is rather challenging with current trial-and-error prompt engineering practices, where developers examine examples in an ad-hoc fashion and iterate their prompts only when they observe outputs that violate their expectations [22].
Second, some requirements may be violated less frequently ($p_{\text{violated}}$), and thus less likely to be discovered via observing violations. Yet they can still be critical, such as high-stakes safety requirements for trip advisory – e.g., “no dangerous activities suggested.” In other cases, violations may be harder to recognize, with a lower probability of being noticed ($p_{\text{noticed}}$), as in “ensuring correct program execution in code explanations.” Moreover, when developers inspect LLM outputs, their assessments are biased by prior knowledge and subjective interpretation, which can lead them to overlook certain types of requirements [38].
Considering these factors, an unspecified requirement may require significant efforts or luck to be discovered with the current practice. Quantitatively, a developer will need to look at $N=\log(1-p_{s})/\log(1-p_{\text{relevant}}\cdot p_{\text{violated}}\cdot p_{\text{noticed}})$ examples to discover the requirement with probability $p_{s}$ (e.g., 0.95), assuming independent Bernoulli trials. For example, a conditional requirement that is relevant to 20% of examples, violated 50% of the time, and perfectly noticeable will require inspecting more than 29 examples to be detected with 95% probability (Figure 4). This will be an excessive workload for a human to complete on their own for every single model update or prompt change, assuming they have access to a diverse dataset, if at all. To overcome these challenges, we envision a more systematic requirements discovery approach in Section 5.
@@ 3.5 LLMs struggle with following many requirements at the same time
Setups. We generated additional prompts containing different numbers of requirements ($N$=1, 5, 10, and 19), following the same method described in Section 3.1. We study their behaviors on two models, Llama-3.3-70B-Instruct and gpt-4o-2024-08-06, and calculate the average accuracy on $N$ specified and $20-N$ unspecified requirements for each prompt.
Results. First, we found that LLMs are mostly able to follow specified requirements individually, with an average of 98.7% accuracy. This can be thought of as an approximate upper bound on performance, assuming each requirement could be stated in isolation without conflict. Next, we found LLMs’ average accuracy starts to drop with more requirements specified (Figure 5): Specifying 19 requirements together yields only an 85.0% average accuracy for gpt-4o. Smaller models like Llama-3.3-70B-Instruct struggle even more, with only 79.7% average accuracy.
@@
3.6 Limitations
While our experiments are conducted on a relatively small number of requirements (n=60) and synthetic prompts (n=240), this setup allows us to curate high-quality, human-validated, realistic requirement-validator pairs and study underspecification systematically. While we rely on LLM-as-a-judge to scale up the evaluation, we manually validate their reliability and ensure a small error rate. To scale up the analysis, future work will need to curate a larger set of realistic requirements and validators – note that high-quality validators usually require manual validation [35]. They will also need more efficient ways (our experiments involved over 1.5 million LLM calls, see Appendix A.8) to produce fine-grained requirement-specific evaluation results while keeping the results reliable.
@@ 4 Requirements-Aware Prompt Optimization
How can we improve LLMs’ performance despite their limited capabilities to follow complex instructions? Inspired by recent work on automatically improving prompts [18], we test whether existing optimizers can already mitigate this issue, and found that they provide inconsistent improvements. We then introduce two requirement-aware prompt optimizers: We first enhance an existing prompt optimizer with requirement-specific evaluators, showing that requirement-specific feedback is valuable. Inspired by our analysis in Section 3, we also explore whether we can optimize what requirements to specify in the prompts, removing ones that are distracting or followed by default. We propose a simple Bayesian prompt optimizer to efficiently search for a good requirement combination.
We first explore whether we can improve prompts with existing prompt optimizers. We use two off-the-shelf prompt optimizers here: (1) OpenAI’s prompt optimizer [29]: This is a “static” prompt optimizer that takes in a prompt and tries to improve it without looking at any model execution feedback. (2) DSPy’s COPRO optimizer [18]: This represents a “dynamic” prompt optimizer that iteratively proposes and explores new prompts and finds ones with higher performance. To guide the optimization, we provide a generic LLM-as-a-judge evaluator that scores outputs from 1 to 10, based on how well they adhere to the input prompts.
@@ 4.1 Existing prompt optimizers do not improve performance consistently
Setups. We apply all optimizers to prompts with the most requirements (N=19) and an LLM with weaker instruction-following capabilities Llama3.3-70b-instruct. We perform prompt optimization on the train split (n=30) of each task dataset, and evaluate the results on the test split. All dynamic prompt optimizers are given a budget of 9 prompts to explore. We measure average requirement accuracy and the number of tokens used in the prompt.
@@ 4.2 Designing requirement-aware prompt optimizers
Bayesian. We then explore optimizing what requirements to specify in a prompt. This is from our observations that many requirements are actually redundant, followed by default, and do not need to be specified (Section 3). To explore an exponential number of requirement combinations, we propose a simple Bayesian prompt optimizer: We view each requirement as a hyperparameter that takes binary values (specified vs. unspecified), and the goal is to select a subset that maximizes the model’s performance on a given task. Formally, let $r=(r_{1},r_{2},\ldots,r_{n})\in\{0,1\}^{n}$ represent the binary configuration of $n$ possible requirements, and let $f(r)$ denote the model’s performance under configuration $r$. The optimizer aims to solve the objective: $r^{*}=\arg\max_{r\in\{0,1\}^{n}}f(r)$.
The performance $f(r)$ here can be defined as any aggregated metric over all requirements that developers care about – we use average requirement accuracy for our experiment. Our prompt optimizer leverages a classic Bayesian optimization algorithm, Tree-structured Parzen Estimator [5], to efficiently find a good $r$ in a small number of trials.
Results. We found that both requirement-aware optimizers can consistently improve prompt performance (Table 2). The simple Bayesian optimizer improves prompt performance by 3.8%, while reducing token usage by 41 - 45%. Pairing requirement-specific validators with existing optimizers, we found COPRO-R is able to further improve performance by 5.8%.
Inspecting prompts produced by COPRO-R (Figure 15), we found them tend to reorder requirements in a more logical structure, merge related requirements together, and sometimes drop requirements from the list. These together produce a better way to specify a longer list of requirements. In contrast, the Bayesian optimizer is able to improve performance simply by deciding what subset of requirements to specify, which is complementary to general-purpose prompt optimizers.
@@ 5 Towards Managing Prompt Underspecification
While requirement-aware prompt optimizers are effective, their success depends on a more rigorous, end-to-end process for building LLM-powered applications. Drawing on insights from software engineering [36], we outline a structured approach to managing underspecification – where prompt developers can discover key task requirements, curate them to reflect real needs, and continuously monitor for drift.
@@
Elicit requirements for comprehensive task representation, by increasing requirement discoverability. As emphasized in the software engineering literature [40], requirement elicitation is foundational for application development. However, as discussed in Section 3.4, identifying task-specific requirements remains nontrivial. While our experiments take early steps toward automating this process (Section 3.1), the broader challenge is to maximize requirement discoverability. This involves increasing $p_{\text{relevant}}$, $p_{\text{noticed}}$, and $p_{\text{violated}}$. To do so, future work may explore using synthetic data generation [47] to surface edge cases or probe specific requirements. We may also explore more traditional requirement engineering approaches, including top-down brainstorming [42], bottom-up analysis that discover requirements from data [44], or safety engineering approaches that use structured processes to anticipate problems before they occur [15].
@@
Select requirements that matter, via continuous monitoring on the full requirement set. LLM behaviors are constantly drifting over time [10] and they have different capabilities to follow requirements that are specified and guess ones that are not (Section 3.2). Deciding what requirements to specify and how to specify them (Section 4), therefore, is largely model-dependent. To support model migrations, we recommend background validation of all tracked requirements. This allows detection of drift and allows developers or optimizer to re-select which requirements to explicitly specify when switching models, from a stable overarching set. This involves always running validations whenever a new prompt or model is used, and alerting developers when significant changes happen to certain individual requirements. Alternatively, instead of single prompts, we could optimize workflow or multi-agent systems [49] where requirements are distributed across sub-modules, enabling more local monitoring and optimization.
@@
Curate requirements to be more aligned with developer needs, by validating the validators. Both optimization and monitoring rely on robust requirement validators. While validators can be separately tuned (e.g., we assessed LLM-as-judge reliability manually), a more rigorous approach is to close the loop between requirements and their validators. If a validator gives unstable (e.g., low-confidence or inconsistent) outputs—especially when compared to human annotations – it likely signals ambiguity or misalignment in the requirement itself. To do so, future work can invest into meta evaluation and alignment of LLM-as-judge [35], strategically have developers review (intentionally curated) validation results, and develop strategies that can refine the requirements. This could include expressing requirements in different representations (use few-shot demonstrations instead of natural language descriptions [18]), decomposing compound goals, or simplifying unrealistic ones.
@@ 6 Related Work
Instruction following capabilities of LLMs. Much research has investigated LLMs’ instruction-following capabilities, from building datasets [26, 50, 31], training better models [33, 30], to optimizing for task-specific instruction-following capabilities [47]. Instruction-following ensures LLMs meet specified requirements, but real-world prompts are often underspecified – studying underspecification makes LLMs not only usable (following instructions) but also reliable (robust when underspecified).
Empirical analysis of LLM behaviors. Lots of work has empirically analyzed the behaviors of LLMs: They found that LLMs are sensitive to prompt design [34, 9], that different LLMs exhibit different behaviors [13, 37], and that LLM updates can often trigger unexpected performance regression [10, 24]. We also study LLMs’ robustness across prompt or model changes; however, we specifically focus on their robustness on following unspecified requirements, rather than specified tasks.
Ambiguity resolution when interacting with LLMs. Ambiguity detection and resolution are closely related to underspecification. For interactive applications, asking clarifying questions can be used as a fallback mechanism to resolve where prompts are underspecified [45, 46]. However, studies found that LLMs often struggle to detect ambiguity in user queries [41, 23], and much work has tried to improve LLMs’ abilities to handle ambiguity [20, 19].
@@ 7 Conclusion
Real-world prompts are underspecified – while LLMs can often infer unspecified requirements, their behaviors are less robust. Underspecification is unavoidable but challenging to address: Unspecified requirements are hard to discover with existing practices, and LLMs struggle with prompts with too many requirements. We demonstrate that requirements-aware optimizers can help produce prompts that are easier to follow, and envision that prompt underspecification can be properly managed through systematic discovery, evaluation, and monitoring processes.
Acknowledgments and Disclosure of Funding
This work is supported by the OpenAI Research Credit program, the Amazon AI Research gift fund, and the Gemma Academic Program GCP Credit Award. We thank Maarten Sap for generously providing us with additional computational resources to run the experiments. We thank Xinran Zhao, Vijay Viswanathan, Jessie Mindel, Jenny T. Liang, Nadia Nahar, Manisha Mukherjee, Vasu Vikram, and Anjiang Wei for their feedback on this work.
@@ A.1 Task descriptions
We selected three tasks based on Anthropic’s report measuring AI usage patterns [14]:
@@ A.2 Process of data sampling and cleaning
We re-purposed three existing datasets to run the LLM+Prompts on: Commitpackft [27] for code explanation (MIT license), a subset of UltraChat [12] for trip advisory (MIT License), and Amazon ESCI [32] for product description generation (Apache-2.0 License).
@@ A.3 Process of requirements curation
. We kept all requirements that are curated from existing prompt, as they already approved by some human developers.
@@
For requirements we elicited with two other approaches, we use text-embedding-ada-002 to generate embeddings of each requirement and remove ones with high cosine similarity ($>0.9$) to other existing requirements incrementally. After this step, we curate 38, 39, and 40 requirements for trip-advisory, product-gen, and code-explain tasks respectively.
We then filtered out the requirements that are overly specific (e.g., “The output must explain how the product’s features enhance the karaoke experience for the targeted age group.”), and finally had three independent annotators select the ones they felt useful for the task.
@@
Figure 7: Existing prompts for the three studied tasks.

### v2

@@ 3.3 LLMs are more likely to regress on unspecified requirements when updated
Figure 3: Cumulative distribution of accuracy drop (truncated at 0%). Prompts regress more on unspecified requirements across model updates, with an almost 2x increase compared to specified requirements.
Setups. To analyze model updates, we study six models from gpt-4o and llama-3 model families: Three versions of gpt-4o: 05-13, 08-06, and 11-20 for simulating a hidden drift of model versions, and three versions of Llama-3-70B-Instruct: Llama-3, Llama-3.1, and Llama-3.3, simulating intentional model migration and bigger changes. For each potential update within the same model family, we calculate the accuracy change for each prompt and requirement.
Results. We found that while the majority of model updates result in stable behaviors (<1% changes) or improvement, there are still a significant portion (22.9%) of cases where LLMs regress. Breaking down the regressions, we found that prompts regress more often on unspecified requirements: 5.9% requirements regress more than 20% over model updates when they are unspecified – an almost 2x increase compared to specified requirements (Figure 3). This is true even for small hidden model updates, – e.g., updating gpt-4o from 05-13 to 08-06 makes it 48% less likely to “produce skimmable outputs,” and 14% less likely to “mention customer support information” on average.
Implications. LLM updates are more likely to improve on specified requirements (through stronger instruction-following) but can hurt unspecified requirements (with different default behaviors). Regressions of unspecified requirements are both more frequent and far harder to detect. This makes it necessary to regularly evaluate and monitor known unspecified requirements (Section 5). Existing practices of manual inspection (“vibe check”) of a few examples will not be sustainable.
@@ 3.4 LLMs struggle with following many requirements at the same time
Setups. We generated additional prompts containing different numbers of requirements (NN=1, 5, 10, and 19), following the same method described in Section 3.1. We study their behaviors on two models, Llama-3.3-70B-Instruct and gpt-4o-2024-08-06, and calculate the average accuracy on $N$ specified and 20N20-N unspecified requirements for each prompt.
Results. First, we found that LLMs are mostly able to follow specified requirements individually, with an average of 98.7% accuracy. This can be thought of as an approximate upper bound on performance, assuming each requirement could be stated in isolation without interference. Next, we found LLMs’ average accuracy starts to drop with more requirements specified (Figure 4): Specifying 19 requirements together yields only an 85.0% average accuracy for gpt-4o. Smaller models like Llama-3.3-70B-Instruct struggle even more, with only 79.7% average accuracy.
@@ 4 Requirements-Aware Prompt Optimization
How can we improve LLMs’ performance despite their limited capabilities to follow complex instructions? Inspired by recent trend on automatically improving prompts Khattab et al. (2023), we test whether existing optimizers can already mitigate this issue, and found that they provide inconsistent improvements. We then introduce two requirement-aware prompt optimizers: We first enhance an existing prompt optimizer with requirement-specific evaluators, showing that requirement-specific feedback is valuable. Inspired by our analysis in Section 3, we also explore whether we can optimize what requirements to specify in the prompts, removing ones that are distracting or followed by default. We propose a simple Bayesian prompt optimizer to efficiently search for a good requirement combination.
@@ 4.1 Existing prompt optimizers do not improve performance consistently
We first explore whether we can improve prompts with existing prompt optimizers. We use two off-the-shelf prompt optimizers here: (1) OpenAI’s prompt optimizer OpenAI (2025): This is a “static” prompt optimizer that takes in a prompt and tries to improve it without looking at any model execution feedback. (2) DSPy’s COPRO optimizer Khattab et al. (2023): This represents a “dynamic” prompt optimizer that iteratively proposes and explores new prompts and finds ones with higher performance. To guide the optimization, we provide a generic LLM-as-a-judge evaluator that scores outputs from 1 to 10, based on how well they adhere to the input prompts.
Setups. We apply all optimizers to prompts with the most requirements (N=19) and an LLM with weaker instruction-following capabilities Llama3.3-70b-instruct. We perform prompt optimization on the train split (n=30) of each task dataset, and evaluate the results on the test split. All dynamic prompt optimizers are given a budget of 9 prompts to explore. We measure average requirement accuracy and prompt token usage.
@@ 4.2 Designing requirement-aware prompt optimizers
Bayesian. We then explore optimizing what requirements to specify in a prompt. This is from our observations that many requirements are actually redundant, followed by default, and do not need to be specified (Section 3.2). To explore an exponential number of requirement combinations, we propose a simple Bayesian prompt optimizer: We model each requirement as a binary hyperparameter (specified vs. unspecified), with a configuration $r=(r_{1},r_{2},\ldots,r_{n})\in\{0,1\}^{n}$ indicating what requirements to specify. Given a performance function $f(r)$, the optimizer aims to solve the objective: $r^{*}=\arg\max_{r\in\{0,1\}^{n}}f(r)$.
We define $f(r)$ as average requirement accuracy for our experiment, and leverage a classic Bayesian optimization algorithm, Tree-structured Parzen Estimator Bergstra et al. (2011), to efficiently find a good $r$ in a small number of trials.
Results. We found that both requirement-aware optimizers can consistently improve prompt performance (Table 2). The simple Bayesian optimizer improves prompt performance by 3.8%, while reducing token usage by 41 - 45%. Inspecting the produced prompts, we found that global, format, and developer-written requirements are more likely to be dropped (56.5%, 77.0%, 57.9% respectively vs. 52.8% avg.), which aligns well with our observations in Section 3.2 that these requirements are more likely to be guessed by an LLM.
Pairing requirement-specific validators with existing optimizers, we found COPRO-R can also improve performance by 5.8%. Inspecting the produced prompts (Figure 16), we found they reorder requirements in a more logical structure, merge related requirements together, and sometimes drop requirements from the list. These together produce a better way to specify a longer list of requirements. In contrast, the Bayesian optimizer offers a complementary strength to improve performance by deciding what subset of requirements to specify.
@@ 5 Towards Managing Prompt Underspecification
While requirement-aware prompt optimizers are effective, their success depends on a more rigorous, end-to-end process for building LLM-powered applications. Drawing on insights from software engineering Sommerville (2015), we outline a structured approach to managing underspecification – where prompt developers can discover key task requirements, curate them to reflect real needs, and continuously monitor for drift.
@@
Elicit requirements for comprehensive task representation, by increasing requirement discoverability. As emphasized in the software engineering literature Van Lamsweerde (2009), requirement elicitation is foundational for application development. However, as discussed in Section 2, identifying task-specific requirements remains nontrivial. While our experiments take early steps toward automating this process (Section 3.1), the broader challenge is to maximize requirement discoverability. Future work may explore using synthetic data generation Zhao et al. (2024) to surface edge cases or probe specific requirements. We may also explore more traditional requirement engineering approaches, including top-down brainstorming Yang et al. (2023), bottom-up data-driven analysis Zeng et al. (2025), or structrued safety engineering approaches that anticipate problems before they occur Hong et al. (2025).
@@
Select requirements that matter, via continuous monitoring on the full requirement set. LLM behaviors drift over time Chen et al. (2024) and they have different capabilities to follow requirements that are specified and guess ones that are not (Section 3.2). Deciding what requirements to specify and how to specify them (Section 4), therefore, is largely model-dependent. To support model migrations, we recommend background validation of all tracked requirements. This allows detection of drift and allows developers or optimizers to re-select which requirements to explicitly specify when switching models, from a stable overarching set. This requires running validations for each new prompt or model and alerting developers of significant changes. Alternatively, instead of single prompts, we could optimize workflow or multi-agent systems Zhou et al. (2025) where requirements are distributed across sub-modules, enabling more local monitoring and optimization.
@@
Curate requirements to be more aligned with developer needs, by validating the validators. Both optimization and monitoring rely on robust requirement validators. While validators can be separately tuned (e.g., we performed manual validation), a more rigorous approach is to close the loop between requirements and their validators. If a validator gives unstable (e.g., low-confidence or inconsistent) outputs – especially when compared to human annotations – it likely signals ambiguity or misalignment in the requirement itself. To do so, future work can invest in meta evaluation and alignment of LLM-as-judge Shankar et al. (2024), strategically have developers review (intentionally curated) validation results, and develop strategies that can refine the requirements.
@@ 6 Related Work
Instruction following capabilities of LLMs. Much research has investigated LLMs’ instruction-following capabilities, from building datasets Zhou et al. (2023); Qin et al. (2024), training better models Sanh et al. (2021); Ouyang et al. (2022), to optimizing for task-specific instruction-following capabilities Zhao et al. (2024). Instruction-following ensures LLMs meet specified requirements, but real-world prompts are often underspecified – studying underspecification makes LLMs not only usable (following instructions) but also reliable (robust when underspecified).
Empirical analysis of LLM behaviors. Lots of work has empirically analyzed the behaviors of LLMs: They found that LLMs are sensitive to prompt design Sclar et al. (2023); Cao et al. (2024), that different LLMs exhibit different behaviors Dunlap et al. (2024); Sun et al. (2025), and that LLM updates can often trigger unexpected performance regression Chen et al. (2024); Ma et al. (2024). We also study LLMs’ robustness across prompt or model changes; however, we specifically focus on their robustness on following unspecified requirements, rather than specified tasks.
Ambiguity resolution when interacting with LLMs. Ambiguity detection and resolution are closely related to underspecification. For interactive applications, asking clarifying questions can be used as a fallback mechanism to resolve where prompts are underspecified Zhang and Choi (2023); Zhang et al. (2024). However, studies found that LLMs often struggle to detect ambiguity in user queries Vijayvargiya et al. (2025); Ma et al. (2025), and much work has tried to improve LLMs’ abilities to handle ambiguity Kim et al. (2024).
@@ 7 Conclusion
Prompt underspecification is common and challenging: LLMs can infer missing requirements but behave inconsistently across prompts and models, while fully specified requirements overwhelm them. We show that requirements-aware optimizers make prompts easier to follow and argue that systematic discovery, evaluation, and monitoring can help effectively manage underspecification.
Limitations
Scale of experiments. Our experiments are conducted on a relatively small number of requirements (n=60) and synthetic prompts (n=240). This setup allows us to curate high-quality, human-validated, realistic requirement–validator pairs and study underspecification systematically, but it limits the breadth of our analysis. To scale up, future work will need to curate a larger set of realistic requirements and validators – note that high-quality validators usually require manual validation Shankar et al. (2024). They will also need more efficient ways (our experiments involved over 1.5 million LLM calls, see Appendix A.8) to produce fine-grained requirement-specific evaluation results while keeping the results reliable.
Evaluation methodology. We rely on LLM validators to scale up the evaluation. While we manually validate their reliability and ensure a small error rate, LLM validators can exhibit preference biases toward models from the same family. This, however, does not affect our main results, which do not compare across model families (e.g., “model updates regress more on unspecified requirements”).
Connection to practices. Our primary goal in this work is to first establish an empirical understanding of the underspecification problem. Accordingly, we do not include a user study or direct integration with developer workflows. This raises the possibility that the observed patterns may not transfer cleanly into practice. Future work should evaluate how well these findings generalize when embedded in real-world processes, constraints, and decision-making contexts of prompt development.
@@ A.1 Task descriptions
We selected three tasks based on Anthropic’s report measuring AI usage patterns Handa et al. (2025):
@@ A.2 Process of data sampling and cleaning
We re-purposed three existing datasets to run the LLM+Prompts on: Commitpackft Muennighoff et al. (2023) for code explanation (MIT license), a subset of UltraChat Ding et al. (2023) for trip advisory (MIT License), and Amazon ESCI Reddy et al. (2022) for product description generation (Apache-2.0 License).
@@ A.3 Process of requirements curation
(Figure 7). We kept all requirements that are curated from existing prompt, as they already approved by some human developers.
@@
Instruction: Review the provided list of requirements and select up to 10 that you believe are the most important for the task.
For each selected requirement, include a brief justification explaining why it is important.
Task description: Explain the code snippet.
Sample inputs:
```python
from sqlalchemy.exc import IntegrityError
from ggrc import db
from ggrc import models
from integration.ggrc import TestCase
class TestCAD(TestCase):
"""Tests for basic functionality of cad model."""
def test_setting_reserved_words(self):
"""Test setting any of the existing attribute names."""
with self.assertRaises(ValueError):
cad = models.CustomAttributeDefinition()
cad.definition_type = "Section"
cad.title = "title"
with self.assertRaises(ValueError):
cad = models.CustomAttributeDefinition()
cad.title = "title"
cad.definition_type = "Section"
Figure 6: Sample annotator instruction during requirement curation.
For requirements we elicited with two other approaches, we use text-embedding-ada-002 to generate embeddings of each requirement and remove ones with high cosine similarity (>0.9>0.9) to other existing requirements incrementally. After this step, we curate 38, 39, and 40 requirements for trip-advisory, product-gen, and code-explain tasks respectively.
We then filtered out the requirements that are overly specific (e.g., “The output must explain how the product’s features enhance the karaoke experience for the targeted age group.”), and finally had three independent annotators select important requirements. We recruited the annotators from our contacts and ensured that they all have sufficient background knowledge for the task they need to annotate (e.g., all three annotators are in a computer science PhD program for the code explanation task). The annotators have access to a short annotation rubric, including an instruction that asks them to mark up important requirements for the task and add justifications when needed, a task description, and example inputs to ground their annotations (Figure 6). The annotation process takes up to 20 minutes, and the annotators are not compensated for their time. We verbally communicated to the annotators that their selected requirements will be used for later experiments.
@@
Figure 7: Existing prompts for the three studied tasks. We extracted a subset of requirements from these prompts and constructed a set of synthetic prompts with the prompt templates in Appendix A.7